# Asynchronous Message-Passing and Zeroth-Order Optimization Based Distributed Learning with a Use-Case in Resource Allocation in Communication Networks

### v2

@@ VIII Numerical Experiments
We consider $m$ agents within the communication network. Each agent $i$ is randomly located within the area $l(i)\in[-m,m]^{2}$. The corresponding receiver $r(i)$ is then randomly located within the area $l(r(i))\in[l(i)-\nicefrac{{m}}{{4}},l(i)+\nicefrac{{m}}{{4}}]$. The considered channel coefficients $h_{r(j)i}$ consist of both large-scale fading, which includes the pathloss gain $h^{p}_{r(j)i}={\mathopen{}\left\|{l(i)-l(r(j))}\right\|\mathclose{}}^{-2.2},\forall i,j\in[m]$, and small-scale fading represented by $h^{f}_{r(j)i}\sim\mathcal{C}\mathcal{N}(0,1)$, with a circularly-symmetric complex normal distribution, hence $h_{r(j)i}=h^{p}_{r(j)i}h^{f}_{r(j)i}$. Transmitter-receiver pairs that are far from each other may have negligible channel gain, and hence, we exclude them from the policy inputs, defined in the next paragraph. For this, we set $\eta=0.01$ and use the modified channel $\tilde{\bm{H}}\in\mathbb{C}^{|\mathcal{R}|\times|\mathcal{T}|}$ whose elements are $\tilde{h}_{r(j)i}=h_{r(j)i}$ if $|h_{r(j)i}|\geq\eta$, and $\tilde{h}_{r(j)i}=0$ otherwise.
@@
Centralized learning is considered the benchmark, where individual policies are trained synchronously on a server. It is noted that this scenario can also be considered as decentralized learning where the underlying graph $\mathcal{G}$ is fully connected, there is no communication delay, and the agents are synchronous, i.e., $T^{i}_{tx}=T^{i}_{u}=T^{i}_{q}=\mathbb{N}$. The experiments are repeated 5 times with different random seeds and the average is reported.
@@ Results
In the first experiment, we assess the performance of the proposed learning approach under varying probabilities $p\coloneqq p_{u}=p_{tx}$, with $m=24$ agents. This parameter, along with the number of neighboring agents $N$, directly impacts the maximum delay $D^{\max}$ in VI.2 which determines the maximum age of the stored quantities $\bar{r}_{i}$ in each agent’s memory. A lower probability $p$ results in a longer propagation time for the quantities $\bar{r}_{i}$ throughout the network (cf. Section VII-B for further explanations). The observations in Fig. 2, where the performance is reported versus iteration number $t$, are aligned with the theoretical findings in VI.2, as increasing $D^{\max}$ leads to slower convergence. In addition, convergence is comparable with centralized learning, even with probabilities as low as $p=0.25$.
Figure 4: Convergence of the proposed asynchronous distributed learning approach with different numbers of agents $m$. Each agent communicates with $N$ neighboring agents and participates in training with probability $p$.. The performance is normalized by the maximum achieved sum rate for each $m$. The comparison is against centralized learning with fully connected agents ($N=m$) in synchronous mode and with no communication delay ($p=1$).
@@ IX Conclusion
In this paper, we have investigated the problem of distributed learning where agents collaborate on a common task while they also have distinct individual tasks. In this context, the performance of each agent is influenced by other agents through interactions. To optimize their local parameters, agents utilize zeroth-order oracles and exchange their local queries through a message-passing architecture, subject to random but bounded communication delays. Importantly, the shared quantities are scalar, ensuring efficient communication bandwidth and preserving agents’ privacy. Moreover, the agents engage in distributed learning in an asynchronous manner. We have also established convergence analyses for smooth nonconvex problems, achieving the same convergence rate as existing learning approaches with synchronous updates. According to the results, the stepsize needs to be decreased for larger delays.
As a use-case, the relevant problem of DL-based RA in communication networks is addressed, where agents, acting as transmitters, collaborate in training their local DNNs using the proposed learning approach to maximize the expected global reward in the communication network. An interesting direction for future research is to extend the proposed learning approach to nonsmooth settings and investigate its application in other learning tasks.
@@ Appendix A Preliminaries
$f_{\ell}$ local reward function $T_{q}^{\ell}$ set of query instants
@@
$F_{\ell}$ local reward value $\mathcal{B}_{\ell},\tilde{\mathcal{B}}_{\ell}$ local memories
@@
$\bar{r}_{\ell}$ shared quantities $D^{\rm max}$ maximum delay time
$\mu$ smoothing parameter $\bm{\xi}$ noise/sample vector
@@
Given a Lipschitz continuous and $L_{g}$-smooth function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the following holds for all $\bm{\theta}^{\prime},\bm{\theta}\in\mathbb{R}^{n}$:
@@
Proposition A.2 (Lipschitz continuity and differentiability of $\tilde{f}_{\ell}$).
@@
Proposition A.3 (Lipschitz differentiability of ${f}^{\mu}_{\ell}$ and $\tilde{f}^{\mu}_{\ell}$).
@@
$=\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{\sum_{i\in\mathcal{N}^{\ell}_{rx}}{\mathopen{}\left(\bar{r}_{i}^{\mathcal{T}}\bm{u}_{\ell}^{\mathcal{T}}+\frac{1}{m}{\nabla}_{\ell}{f}^{\mu}_{i}(\bm{\theta}^{\mathcal{T}})\right)\mathclose{}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$
@@
$\mathbb{E}\bigg{[}\bigg{\|}\frac{\frac{1}{B}\sum_{q=1}^{B}{\mathopen{}\left(F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})\right)\mathclose{}}}{\mu}\bm{u}_{\ell}^{\mathcal{T}}$
@@
$\mathbb{E}_{\bm{\xi},\bm{u}}\bigg{[}\bigg{\|}\frac{F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})}{\mu}\bm{u}_{\ell}^{\mathcal{T}}+{\nabla}_{\ell}{f}^{\mu}_{i}(\bm{\theta}^{\mathcal{T}})\bigg{\|}^{2}\bigg{]}$
@@
$\mathbb{E}_{\bm{\xi}}{\mathopen{}\left[{\mathopen{}\left\|{F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{\prime^{q}})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}\mathbb{E}_{\bm{u}}{\mathopen{}\left[{\mathopen{}\left\|{\bm{u}_{\ell}^{\mathcal{T}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$ (23)
where in the first inequality Jensen’s inequality is applied with taking into account $|\mathcal{N}^{\ell}_{rx}|\leq m$. In the second equality, $\bar{r}_{i}$ is expanded using (4), where the batch size is $B$. Also, $F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})$ is added and subtracted. For the last inequality, Jensen’s inequality is applied for both terms, while for the first term Item 2 is also considered. We continue to bound the first term in the last inequality:
$\mathbb{E}_{\bm{\xi},\bm{u}}{\mathopen{}\left[{\mathopen{}\left\|{\frac{F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})}{\mu}\bm{u}_{\ell}^{\mathcal{T}}+{\nabla}_{\ell}{f}^{\mu}_{i}(\bm{\theta}^{\mathcal{T}})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$
$\leq\mathbb{E}_{\bm{\xi},\bm{u}}{\mathopen{}\left[{\mathopen{}\left\|{\frac{F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})}{\mu}\bm{u}_{\ell}^{\mathcal{T}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$
@@ Appendix B Convergence study
We make use of the descent lemma (cf. A.1) on the $L_{\mu}$-smooth function ${f}^{\mu}(\bm{\theta}^{t})$ with an expectation conditioned on the filtration $\mathcal{F}^{t}$, defined in (21). Note that as conditioned on $\mathcal{F}^{t}$, all the iterates up to and including $\bm{\theta}^{t}$ are deterministic:
@@
$\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{{\nabla}_{i}\tilde{f}_{i}(\Theta^{t}_{i})-{\nabla}_{i}\tilde{f}^{\mu}_{i}(\Theta^{t}_{i})+{\nabla}_{i}\tilde{f}^{\mu}_{i}(\Theta^{t}_{i})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$ (30)
$\leq\frac{1}{2m}\mu^{2}L^{2}(n+3)^{3}+2\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{{\nabla}_{i}\tilde{f}^{\mu}_{i}(\Theta^{t}_{i})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}},$
@@
where the first and second inequalities are due to Item 3 and Item 2, respectively. Putting the last bound back into (31) results in
@@
where in the first inequality, Young’s inequality and (15), and in the second inequality, $\tilde{L}$-smoothness of $\tilde{f}_{\ell}$ established by A.2 are considered. The bound is also valid if both sides are multiplied by $\alpha^{t}\coloneqq\nicefrac{{\bar{\gamma}^{t}}}{{\sum_{k=0}^{\bar{t}}\bar{\gamma}^{k}}}$ where $\bar{\gamma}^{k}$ is defined in (28) and are summed over $t$:
@@
With the defined stepsize sequence, the first term in the rhs is diminishing wrt $\bar{t}$ due to (32). Similarly, the last term in the rhs is diminishing due to (29). Inequalities (29) and (32) are considered with $\tilde{\sigma}^{2}$ defined in VI.3 here since the assumption that the oracles $F_{\ell}(\cdot,\bm{\bm{\xi}}),\forall\ell\in[N]$ can be queried in any two points holds. As the last term in (33) dominates the first term in the rhs, it can be concluded that the rhs diminishes with a rate of $\mathcal{O}(\mu^{2})+\mathcal{O}(\nicefrac{{1}}{{\sqrt{\bar{t}}}})$, due to (29) and (30). The lhs can also be lowerbounded by the ${}\min\{\cdot\}$ operator, similar to (29), completing the proof. ∎

### v3

@@ VIII Numerical Experiments
We consider $m$ agents within the communication network with the same number of receivers n=m. Each agent $i$ is randomly located within the area $l(i)\in[-m,m]^{2}$. The corresponding receiver $r(i)$ is then randomly located within the area $l(r(i))\in[l(i)-\nicefrac{{m}}{{4}},l(i)+\nicefrac{{m}}{{4}}]$. The considered channel coefficients $h_{r(j)i}$ consist of both large-scale fading, which includes the pathloss gain $h^{p}_{r(j)i}={\mathopen{}\left\|{l(i)-l(r(j))}\right\|\mathclose{}}^{-2.2},\forall i,j\in[m]$, and small-scale fading represented by $h^{f}_{r(j)i}\sim\mathcal{C}\mathcal{N}(0,1)$, with a circularly-symmetric complex normal distribution, hence $h_{r(j)i}=h^{p}_{r(j)i}h^{f}_{r(j)i}$. Transmitter-receiver pairs that are far from each other may have negligible channel gain, and hence, we exclude them from the policy inputs, defined in the next paragraph. For this, we set $\eta=0.01$ and use the modified channel $\tilde{\bm{H}}\in\mathbb{C}^{m\times m}$ whose elements are $\tilde{h}_{ji}=h_{r(j)i}$ if $|h_{r(j)i}|\geq\eta$, and $\tilde{h}_{ji}=0$ otherwise.
@@
Centralized learning is considered the benchmark, where individual policies are trained synchronously on a server. It is noted that this scenario can also be considered as distributed learning where the underlying graph $\mathcal{G}$ is fully connected, there is no communication delay, and the agents are synchronous, i.e., $T^{i}_{tx}=T^{i}_{u}=T^{i}_{q}=\mathbb{N}$. The experiments are repeated 5 times with different random seeds and the average is reported.
@@ Results
In the first experiment, we assess the performance of the proposed learning approach under varying probabilities $p\coloneqq p_{u}=p_{tx}$, with $m=24$ agents. This parameter, along with the number of neighboring agents $N$, directly impacts the maximum delay $D^{\max}$ in VI.2 which determines the maximum age of the stored quantities $\bar{r}_{i}$ in each agent’s memory. A lower probability $p$ results in a longer propagation time for the quantities $\bar{r}_{i}$ throughout the network (refer to Section VII-B for further explanations). The observations in Fig. 2, where the performance is reported versus iteration number $t$, are aligned with the theoretical findings in VI.2, as increasing $D^{\max}$ leads to slower convergence. In addition, convergence is comparable with centralized learning, even with probabilities as low as $p=0.25$.
Figure 4: Convergence of the proposed asynchronous distributed learning approach with different numbers of agents $m$. Each agent communicates with $N$ neighboring agents and participates in training with probability $p$. The performance is normalized by the maximum achieved sum rate for each $m$. The comparison is against centralized learning with fully connected agents ($N=m$) in synchronous mode and with no communication delay ($p=1$).
@@ IX Conclusion
In this paper, we have investigated the problem of distributed learning where agents collaborate on a common task while they also have distinct individual tasks. In this context, the performance of each agent is influenced by other agents through interactions. To optimize their local parameters, agents utilize zeroth-order oracles and exchange their local queries through a message-passing architecture, subject to random but bounded communication delays. Importantly, the shared quantities are scalar, ensuring efficient communication bandwidth usage and preserving agents’ privacy. Moreover, the agents engage in distributed learning in an asynchronous manner. We have also established convergence analyses for smooth nonconvex problems, achieving the same convergence rate as existing learning approaches with synchronous updates. According to the analysis results, the stepsize needs to be decreased for larger delays.
As a use-case, the relevant problem of DL-based RA in communication networks is addressed, where agents, acting as transmitters, collaborate in training their local DNNs using the proposed learning approach to maximize the expected global reward in the communication network. An interesting direction for future research is to extend the proposed learning approach to nonsmooth regularized settings and investigate its application in other learning tasks.
@@ Appendix A Preliminaries
$f_{\ell}$ local cost function $T_{q}^{\ell}$ set of query instants
@@
$F_{\ell}$ local cost value $\mathcal{B}_{\ell},\tilde{\mathcal{B}}_{\ell}$ local memories
@@
$\bar{r}_{\ell}$ local queries $D^{\rm max}$ maximum delay time
$\mu$ smoothing parameter $\bm{\xi}$ sample vector
@@
Given a continuous and $L_{g}$-smooth function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$, the following holds for all $\bm{\theta}^{\prime},\bm{\theta}\in\mathbb{R}^{n}$:
@@
Proposition A.2 (Lipschitz continuous and differentiable $\tilde{f}_{\ell}$).
@@
Proposition A.3 (Lipschitz differentiable $f^{\mu}_{\ell}$ and $\tilde{f}^{\mu}_{\ell}$).
@@
$=\mathbb{E}_{\bm{\xi},\mathbb{U}}{\mathopen{}\left[{\mathopen{}\left\|{\sum_{i\in\mathcal{N}^{\ell}_{rx}}{\mathopen{}\left(\bar{r}_{i}^{\mathcal{T}}\bm{u}_{\ell}^{\mathcal{T}}+\frac{1}{m}{\nabla}_{\ell}f^{\mu}_{i}(\bm{\theta}^{\mathcal{T}})\right)\mathclose{}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$
@@
$\mathbb{E}_{\bm{\xi},\mathbb{U}}\bigg{[}\bigg{\|}\frac{\frac{1}{B}\sum_{q=1}^{B}{\mathopen{}\left(F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})\right)\mathclose{}}}{\mu}\bm{u}_{\ell}^{\mathcal{T}}$
@@
$\mathbb{E}_{\bm{\xi},\mathbb{U}}\bigg{[}\bigg{\|}\frac{F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})}{\mu}\bm{u}_{\ell}^{\mathcal{T}}+{\nabla}_{\ell}f^{\mu}_{i}(\bm{\theta}^{\mathcal{T}})\bigg{\|}^{2}\bigg{]}$
@@
$\mathbb{E}_{\bm{\xi}}{\mathopen{}\left[{\mathopen{}\left\|{F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{\prime^{q}})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}\mathbb{E}_{\mathbb{U}}{\mathopen{}\left[{\mathopen{}\left\|{\bm{u}_{\ell}^{\mathcal{T}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$ (23)
where in the first inequality Jensen’s inequality is applied with taking into account $|\mathcal{N}^{\ell}_{rx}|\leq m$. In the second equality, $\bar{r}_{i}$ is expanded using (4), where the batch size is $B$. Also, $F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})$ is added and subtracted. For the last inequality, Jensen’s inequality is applied for both terms, while for the first term Item 2 is also considered, with $\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{\frac{1}{B}\sum_{q=1}^{B}{\nabla}H(\bm{\theta},\bm{\xi}^{q})-\mathbb{E}{\mathopen{}\left[{\nabla}H(\bm{\theta},\bm{\xi})\mid{{}}\right]\mathclose{}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}=\frac{1}{B^{2}}\sum_{q=1}^{B}\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{{\nabla}H(\bm{\theta},\bm{\xi}^{q})-\mathbb{E}{\mathopen{}\left[{\nabla}H(\bm{\theta},\bm{\xi})\mid{{}}\right]\mathclose{}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$, a function $H$, and sample vectors $\bm{\xi}^{q}$ and $\bm{\xi}$. We continue to bound the first term in the last inequality:
$\mathbb{E}_{\bm{\xi},\mathbb{U}}{\mathopen{}\left[{\mathopen{}\left\|{\frac{F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})}{\mu}\bm{u}_{\ell}^{\mathcal{T}}+{\nabla}_{\ell}f^{\mu}_{i}(\bm{\theta}^{\mathcal{T}})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$
$\leq\mathbb{E}_{\bm{\xi},\mathbb{U}}{\mathopen{}\left[{\mathopen{}\left\|{\frac{F_{i}(\bm{\theta}^{\mathcal{T}}+\mu\bm{u}^{\mathcal{T}},\bm{\xi}^{q})-F_{i}(\bm{\theta}^{\mathcal{T}},\bm{\xi}^{q})}{\mu}\bm{u}_{\ell}^{\mathcal{T}}}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}$
@@ Appendix B Convergence study
We make use of the descent lemma (refer to A.1) on the $L_{\mu}$-smooth function ${f}^{\mu}(\bm{\theta}^{t})$ with an expectation conditioned on the filtration $\mathcal{F}^{t}$, defined in (21). Note that as conditioned on $\mathcal{F}^{t}$, all the iterates up to and including $\bm{\theta}^{t}$ are deterministic:
@@
$\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{{\nabla}_{i}\tilde{f}_{i}(\Theta^{t}_{i})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}}=\mathbb{E}{{\mathopen{}\left\|{{\nabla}_{i}\tilde{f}_{i}(\Theta^{t}_{i})-{\nabla}_{i}\tilde{f}^{\mu}_{i}(\Theta^{t}_{i})+{\nabla}_{i}\tilde{f}^{\mu}_{i}(\Theta^{t}_{i})}\right\|\mathclose{}}^{2}}$
$\leq\frac{1}{2m}\mu^{2}L^{2}(n+3)^{3}+2\mathbb{E}{\mathopen{}\left[{\mathopen{}\left\|{{\nabla}_{i}\tilde{f}^{\mu}_{i}(\Theta^{t}_{i})}\right\|\mathclose{}}^{2}\mid{{}}\right]\mathclose{}},$ (30)
@@
where the first and second inequalities are due to Item 3, with $\tilde{L}\leq\frac{1}{\sqrt{m}}L$ due to A.2, and Item 2, respectively. Putting the last bound back into (31) results in
@@
where in the first inequality, Young’s inequality and (15), and in the second inequality, $\tilde{L}$-smoothness of $\tilde{f}_{\ell}$ established by A.2 are considered. The bound is also valid if both sides are multiplied by $\alpha^{t}\coloneqq\nicefrac{{\bar{\gamma}^{t}}}{{\sum_{k=0}^{\bar{t}}\bar{\gamma}^{k}}}$, where $\bar{\gamma}^{k}$ is defined in (28), and summed over $t$:
@@
With the defined stepsize sequence, the first term in the rhs is diminishing wrt $\bar{t}$ due to (32). Moreover, as the second term dominates the first term on the rhs, it can be concluded that the rhs diminishes at a rate of $\mathcal{O}(\mu^{2})+\mathcal{O}(\nicefrac{{1}}{{\sqrt{\bar{t}}}})$, due to (29) and (B). It is highlighted that inequalities (29) and (32) are considered here with $\tilde{\sigma}^{2}$ defined in VI.3, since the assumption that the oracles $F_{\ell}(\cdot,\bm{\xi}),\forall\ell\in[N]$, can be queried in any two points holds. The lhs in (33) can also be lowerbounded by the ${}\min\{\cdot\}$ operator, similar to (29), thus completing the proof. ∎